OpenStudy (loganator):

Please Help! Algebra 1 Question Medal and Fan! ;)

OpenStudy (sagewilson):

What is your question.

OpenStudy (loganator):

There it is

OpenStudy (loganator):

So hard! Don't understand how to solve

OpenStudy (loganator):

Thank you for taking a look, I'm in deep trouble if I don't get this assignment in and I don't know where to start

jhonyy9 (jhonyy9):

A. y=2^-x y=4^x+3 by start y=y so than you need make the right side equal too and see what will get for x do you can solve it for x ?

OpenStudy (loganator):

I think I know part A, which would be: The solution would be where the lines intersect because the point where the intersect would represent a point that provides solutions for both sets of equations. however Part B and Part C I am lost

jhonyy9 (jhonyy9):

2^-x = 4^x+3 do you have any idea for how you solve it ?

OpenStudy (loganator):

No, I haven't done that lesson in a while, I would assume that you get the 2 x variables on one side.

jhonyy9 (jhonyy9):

because on the left side there is 2 with an exponent so how can you rewrite the right side to be there 2 with an exponent same to the left side ?

OpenStudy (loganator):

You multiply the left side by 2 and the 2 variables cancel out?

jhonyy9 (jhonyy9):

please concentrate just to the right side now please how can you rewriting the 4^x+3 so that being 2 on an exponent ?

OpenStudy (loganator):

I don't know how to make the exponent 2, I can turn the Base, which is 4, in to 2 simply by multiplying it by one half.

jhonyy9 (jhonyy9):

no for example 9^(x+2) = 3^2(x+2) are these equales ? so than yes please justifie me why ?

OpenStudy (loganator):

because 3 to the second power is 9

jhonyy9 (jhonyy9):

exactly so just you need using these property in case of your exercise too - ok. ?

OpenStudy (loganator):

How does this help me solve the problem

OpenStudy (loganator):

I just don't see how things fit, sometimes I have trouble with that.

jhonyy9 (jhonyy9):

4^x+3 = 2^?

jhonyy9 (jhonyy9):

use exactly what i have wrote to you above

jhonyy9 (jhonyy9):

hope you know that 4=2^2

jhonyy9 (jhonyy9):

so than 4^x+3 = 2^2(x+3) yes ?

OpenStudy (loganator):

Does this apply to part A?

jhonyy9 (jhonyy9):

so than this will be on the right side and now what is on the left side ?

OpenStudy (loganator):

How does this relate to part A?

OpenStudy (loganator):

Not questioning your help I'm just confused

jhonyy9 (jhonyy9):

but than you check what above you have wrote so you have said me that part A you know it

OpenStudy (loganator):

Well it says to Explain how they are solutions when they intersect, it seems like they're looking for an explanation.

jhonyy9 (jhonyy9):

ok. so than you try graphing two lines for eaxmple with coordinates (3,0) and (5,0) i said 0 just for you can see it more easy - being the point origo - of the axe x and y - ok.? so than you graph these lines - looke it please beacuse are y coordinates the same 0 and 0 so this mean that in this point these lines will intersect sure sorry but i think that in an other example i dont explain to you more easy do you understand it now ?

jhonyy9 (jhonyy9):

so from this result that where y=y these lines will intersect sure - this is for part A for part B - check please what i have clarified with you above

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (loganator):

Yes

OpenStudy (loganator):

@jhonyy9

OpenStudy (loganator):

@jhonyy9

OpenStudy (loganator):

Omg, jerk

jhonyy9 (jhonyy9):

do you understand what i have wrote above ?

jhonyy9 (jhonyy9):

for part A ?